OpenStudy (anonymous):

please help will give medal. Simplify the expression 9+3x2

OpenStudy (nirmalnema):

3(3+x^2)

OpenStudy (mokeira):

\[9+3x ^{2}=3^{2}+3x ^{2}\]

OpenStudy (anonymous):

here are the answers a 15 b 21 c 24 d 48

OpenStudy (nirmalnema):

ohh ok. 9+(3*2)=9+6 =15